Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Pushing Probability to Punishing Exhaustion

Expert replies
by dtweah » Wed May 13, 2009 5:02 pm
A and B play a series of games; the winner of each game has probability 0.8 of winning the next game. If A wins the first game, what is the probability A will win the third game?

(a) .68
(b) .76
(c) .72
(d) .60
(e) .64
Join the discussion
Source: — Problem Solving |

by ontherocks27 » Wed May 13, 2009 11:29 pm
There are two ways in which A can win the 3rd game after winning the first game

1st way: A wins 2nd game and A wins 3rd game
Probablity = 0.8 * 0.8 = 0.64

or

2nd way: B wins 2nd game and A wins 3rd game
Probablity = (1-0.8) * (0.2) = 0.2*0.2 = 0.04

Overall probability of A winning the 3rd game = 0.64+0.04=0.68

Hence (A)
Join the discussion

by sanju09 » Thu May 14, 2009 4:54 am
After A wins the first game, A can win the third game in two situations

1. A loses second and wins the third: probability = 0.2 × 0.2 = 0.04. OR
2. A wins second and also the third: probability = 0.8 × 0.8 = 0.64.

Total probability = 0.04 + 0.64 + 0.68. My [spoiler]A[/spoiler].

Explanation: In case 1, A loses second means B wins, and A wins third means B loses. See how the probabilities change as per the condition given. Hope no doubts.
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion